OpenStudy (anonymous):

okay I have 4 but one at a time.... someone unstump me please. 1) Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0, -3).

OpenStudy (anonymous):

My brain gets jumbled too easy so keep it simple please

OpenStudy (anonymous):

this one we do the vertex way if the vertex is \((h,k)\) we know it looks like \[y=a(x-h)^2+k\] simple or no?

OpenStudy (anonymous):

replace \(h\) by 0, \(k\) by -3 and write \[y=a(x-0)^2-3\] i.e. \[y=ax^2-3\] the only number you do not know is \(a\)

OpenStudy (anonymous):

meh... kinda. this has my mind completely lost :/

OpenStudy (anonymous):

OHHH!

OpenStudy (anonymous):

then take a break from it, there is no other way to do it, use the roots, or use the vertex

OpenStudy (anonymous):

I cant. I need to finish it

OpenStudy (anonymous):

please finish this one :)

OpenStudy (anonymous):

Ill try the next one with your guidance

OpenStudy (anonymous):

we know that \[f(1)=0\] so to find \(a\) replace \(x\) by 1, set the result equal to zero, you have \[f(x)=ax^2-4\] \[f(1)=a(1)^2-3=0\] \[a-3=0\] \[a=3\] answer is \[f(x)=3x^2-3\]

OpenStudy (anonymous):

typo on first line, should have been \[f(x)=ax^2-3\]

OpenStudy (anonymous):

I kind of get it (:

OpenStudy (anonymous):

can we do the next one? Guide me?